OpenStudy (anonymous):

the acc of free fall on a planet P is 1\6 of the free fall on earth. the mass of the body on planet P is 30kg. what is the weight on planet P?

OpenStudy (anonymous):

w = mg

OpenStudy (anonymous):

g on earth is 9.8 m/s^2

OpenStudy (anonymous):

On Planet P = g/6 W = 30 * g/6 = 5 * 9.8 = 49 N

OpenStudy (anonymous):

YES its 9.81